OpenStudy (superhelp101):

How many grams of potassium chloride must decompose to produce 69.9 milliliters of potassium metal, if the density of potassium is 1.78 g/mL? Show all steps of your calculation as well as the final answer. KCl → K + Cl2

OpenStudy (superhelp101):

@Abmon98 can u help me?

OpenStudy (somy):

Density= mass/volume you are given density and volume so find Mass

OpenStudy (somy):

convert volume to Liters beforehand

OpenStudy (somy):

ah no don't do it since in unit of density its mL

OpenStudy (somy):

just put the values into the formula and get the mass

OpenStudy (somy):

after you get mass of Potassium use formula mole= mass/Molecular mass and get mole of Potassium

OpenStudy (superhelp101):

but first don't i have to get hte balanced equation? Which is 2K+Cl2=2KCl

OpenStudy (somy):

after you get mole of potassium balance the reaction and look at ratio between K and KCl

OpenStudy (somy):

is it clear till now?

OpenStudy (somy):

afterwards by ratio you will find mole of KCl and as you find the mole use formula mole= mass/Molecular mass and make mass as subject mass= mole* Molecular mass and find mass of KCl

OpenStudy (superhelp101):

more clear. thank you

OpenStudy (somy):

you are welcome :))

OpenStudy (abmon98):

Balance Chemical equation 2KCl-->2K+Cl2 Number of moles=Mass(g)/Molar Mass Density=Mass/Volume Mass=Density*Volume 1.78*69.9=124.422 grams produced. Number of Moles 124.422/39.1=3.18 Look at the mole ratio of KCL to K its 2:2 they share the same number of moles Mass=Number of Moles*Molar Mass 3.18*(39.1*1)+(35.5*1) 3.18*74.6=237 grams

OpenStudy (somy):

awesome job @Abmon98 !! but don't do the whole calculation, @superhelp101 has to do it :)

OpenStudy (superhelp101):

wow that you guys!:-) don't worry i will do the calculation once myself :)

OpenStudy (abmon98):

thank you :) @somy and i wont again sorry @superhelp101

OpenStudy (superhelp101):

you're fine

OpenStudy (somy):

:]